OpenStudy (anonymous):

12x+13y=29 and 13x+12y=21 find (x+y)

OpenStudy (anonymous):

please urgent help needed

OpenStudy (anonymous):

i guess no help for me :(

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

try to subtract the one equation from the another

OpenStudy (anonymous):

-1x+1y=8

OpenStudy (anonymous):

now?

OpenStudy (anonymous):

-x+y=8

OpenStudy (anonymous):

x=y-8

OpenStudy (anonymous):

put this in one of the equation

OpenStudy (anonymous):

12(y-8)+13y=29 12y-96+13y=29 25y=29+96 25y=125 y=125/25 y=5

OpenStudy (anonymous):

x=y-8 x=5-8=-3 then x+y=-3+5=2